# LinFusion: 1 GPU, 1 Minute, 16K Image

### v2

@@ Abstract
Modern diffusion models, particularly those utilizing a Transformer-based UNet for denoising, rely heavily on self-attention operations to manage complex spatial relationships, thus achieving impressive generation performance. However, this existing paradigm faces significant challenges in generating high-resolution visual content due to its quadratic time and memory complexity with respect to the number of spatial tokens. To address this limitation, we aim at a novel linear attention mechanism as an alternative in this paper. Specifically, we begin our exploration from recently introduced models with linear complexity, e.g., Mamba2, RWKV6, Gated Linear Attention, etc, and identify two key features—attention normalization and non-causal inference—that enhance high-resolution visual generation performance. Building on these insights, we introduce a generalized linear attention paradigm, which serves as a low-rank approximation of a wide spectrum of popular linear token mixers. To save the training cost and better leverage pre-trained models, we initialize our models and distill the knowledge from pre-trained StableDiffusion (SD). We find that the distilled model, termed LinFusion, achieves performance on par with or superior to the original SD after only modest training, while significantly reducing time and memory complexity. Extensive experiments on SD-v1.5, SD-v2.1, and SD-XL demonstrate that LinFusion delivers satisfactory zero-shot cross-resolution generation performance, generating high-resolution images like 16K resolution. Moreover, it is highly compatible with pre-trained SD components, such as ControlNet and IP-Adapter, requiring no adaptation efforts. Codes are available here.
@@ 1 Introduction
The success of models like SD can be largely attributed to their robust backbone structures for denoising. From UNet architectures with attention layers Ronneberger et al. (2015); Rombach et al. (2022) to Vision Transformers Peebles & Xie (2023); Bao et al. (2023); Chen et al. (2023); Esser et al. (2024), existing designs rely heavily on self-attention mechanisms to manage complex relationships between spatial tokens. Despite their impressive performance, the quadratic time and memory complexity inherent in self-attention operations poses significant challenges for high-resolution visual generation. For instance, as illustrated in Fig. 2(a), using FP16 precision, SD-v1.5 fails to generate 2048-resolution images on A100, a GPU with 80GB of memory, due to out-of-memory errors, making higher resolutions or larger models even more problematic.
@@
The proposed generalized linear attention module is integrated into the architectures of SD, replacing the original self-attention layers, and the resultant model is termed as Linear-Complexity Diffusion Model, or LinFusion in short. By only training the linear attention modules for 50k iterations in a knowledge distillation framework, LinFusion achieves performance on par with or even superior to the original SD, while significantly reducing time and memory complexity, as shown in Fig. 2. Meanwhile, it delivers satisfactory zero-shot cross-resolution generation performance and can generate images at 16K resolution on a single GPU. It is also compatible with existing components for SD, such as ControlNet Zhang et al. (2023) and IP-Adapter Ye et al. (2023), allowing users to inject additional controls to the proposed LinFusion flexibly without any additional training cost. As shown in Fig. 1, extensive experiments on SD-v1.5, SD-v2.1, and SD-XL validate the effectiveness of the proposed model and method. Our contributions can be summarized as follows:
@@
  • Extensive experiments on SD demonstrate that the proposed LinFusion can achieve even better results than the original SD and exerts satisfactory zero-shot cross-resolution generation performance and compatibility with existing components for SD. To the best of our knowledge, this is the first exploration of linear-complexity token mixers on the SD series model for text-to-image generation.
2 Related Works
2.1 Efficient Diffusion Architectures
To mitigate the efficiency issue triggered by the quadratic time and memory complexity, a series of works, including DiS Fei et al. (2024a), DiM Teng et al. (2024), DiG Zhu et al. (2024), Diffusion-RWKV Fei et al. (2024b), DiffuSSM Yan et al. (2024), and Zigma Hu et al. (2024). These works have successfully adapted recent state space models like Mamba Gu & Dao (2023), RWKV Peng et al. (2023), or Linear Attention Katharopoulos et al. (2020) into diffusion architectures. However, these architectures maintain a causal restriction for diffusion task, processing input spatial tokens one by one, with generated tokens conditioned only on preceding tokens. In contrast, the diffusion task allows models to access all noisy tokens simultaneously, making the causal restriction unnecessary. To address this, we eliminate the causal restriction and introduce a non-causal token mixer specifically designed for the diffusion model.
2.2 Linear-Complexity Token Mixers
@@ 3.2 Overview
Technically, to derive a linear-complexity diffusion backbone, one simple solution is to replace all the self-attention blocks with Mamba2, as shown in Fig. 4(a). We apply bi-directional SSM to ensure that the current position can access information from subsequent positions. Moreover, the self-attention modules in Stable Diffusion do not incorporate gated operations Hochreiter & Schmidhuber (1997); Cho (2014) or RMS-Norm Zhang & Sennrich (2019) as used in Mamba2. As shown in Fig. 4(b), we remove these structures to maintain the consistency and result in a slight improvement in performance. In the following parts of this section, we delve into the issues of applying SSM, the core module in Mamba2, to diffusion models and accordingly introduce the key features in LinFusion: normalization and non-causality in Secs. 3.3 and 3.4 respectively. Finally, in Sec. 3.5, we provide the training objectives to optimize parameters in LinFusion modules.
@@ 3.4 Non-Causal Mamba
Proof.
Given existing conditions, we have:
$=[(c(X_{i})\otimes f(X_{i}))(b(X_{j})\otimes g(X_{j}))^{\top}]_{i,j},$
In practice, we adopt two MLPs to mimic the functionalities of $f^{\prime}$ and $g^{\prime}$. Combining with the normalization operations mentioned in Sec. 3.3, we derive an elegant structure shown in Fig. 4(d).
@@
Proof.
Given existing conditions, we have:
From this point of view, the proposed structure can be deemed as a generalized linear attention and a non-causal form of recent linear-complexity sequential models, including Mamba2 Dao & Gu (2024), RWKV6 Peng et al. (2024), GLA Yang et al. (2023b), etc. In Tab. 1, we provide a summary of the parameterization in recent works for $A_{i}$.
@@ 4.2 Main Results
Ultrahigh-Resolution Generation. As discussed in Huang et al. (2024); He et al. (2024), directly applying diffusion models trained on low resolutions for higher-resolution generation can result in content distortion and duplication. In this paper, we address the challenge by handling the low resolution first, based on which higher-scale images are generated using SDEdit Meng et al. (2021). Note that in this work we aim at a linear-complexity model computationally enabling such ultrahigh-resolution generation as shown in Fig. 1. Dedicated designs are left as future directions.
@@ 4.3 Empirical Extensions
The proposed LinFusion is highly compatible with various components/plugins for SD, such as ControlNet Zhang et al. (2023), IP-Adapter Ye et al. (2023), and LoRA Hu et al. (2022), without any further training or adaptation. We present qualitative results in Fig. 5 and refer readers to the appendix for more results. The following quantitative evaluations indicate comparable performance of LinFusion with the original SD.
@@ 5 Conclusion
This paper introduces a diffusion backbone termed LinFusion for text-to-image generation with linear complexity in the number of pixels. At the heart of LinFusion lies in a generalized linear attention mechanism, distinguished by its normalization-aware and non-causal operations—key aspects overlooked by recent linear-complexity token mixers like Mamba, Mamba2, and GLA. We reveal theoretically that the proposed paradigm serves as a general low-rank approximation for the non-causal variants of recent models. Based on Stable Diffusion (SD), LinFusion modules after knowledge distillation can seamlessly replace self-attention layers in the original model, ensuring that LinFusion is highly compatible to existing components for Stable Diffusion, like ControlNet, IP-Adapter, and LoRA, without any further training effort. Extensive experiments on SD-v1.5, SD-v2.1, and SD-XL demonstrate that the proposed model outperforms existing baselines and achieves performance on par with, or better than, the original SD with significantly reduced computational overhead. On a single GPU, it can accommodate image generation with resolutions up to 16K.
@@
Table 3: Results of ControlNet on the original SD-v1.5 and LinFusion.

### v3

@@ Abstract
Modern diffusion models, particularly those utilizing a Transformer-based UNet for denoising, rely heavily on self-attention operations to manage complex spatial relationships, thus achieving impressive generation performance. However, this existing paradigm faces significant challenges in generating high-resolution visual content due to its quadratic time and memory complexity with respect to the number of spatial tokens. To address this limitation, we aim at a novel linear attention mechanism as an alternative in this paper. Specifically, we begin our exploration from recently introduced models with linear complexity, e.g., Mamba2, RWKV6, Gated Linear Attention, etc, and identify two key features—attention normalization and non-causal inference—that enhance high-resolution visual generation performance. Building on these insights, we introduce a generalized linear attention paradigm, which serves as a low-rank approximation of a wide spectrum of popular linear token mixers. To save the training cost and better leverage pre-trained models, we initialize our models and distill the knowledge from pre-trained StableDiffusion (SD). We find that the distilled model, termed LinFusion, achieves performance on par with or superior to the original SD after only modest training, while significantly reducing time and memory complexity. Extensive experiments on SD-v1.5, SD-v2.1, and SD-XL demonstrate that LinFusion enables satisfactory and efficient zero-shot cross-resolution generation, accommodating ultra-resolution images like 16K on a single GPU. Moreover, it is highly compatible with pre-trained SD components and pipelines, such as ControlNet, IP-Adapter, DemoFusion, DistriFusion, etc, requiring no adaptation efforts. Codes are available here.
@@ 1 Introduction
The success of models like SD can be largely attributed to their robust backbone structures for denoising. From UNet architectures with attention layers Ronneberger et al. (2015); Rombach et al. (2022) to Vision Transformers Peebles & Xie (2023); Bao et al. (2023); Chen et al. (2023); Esser et al. (2024), existing designs rely heavily on self-attention mechanisms to manage complex relationships between spatial tokens. Despite their impressive performance, the quadratic time and memory complexity inherent in self-attention operations poses significant challenges for high-resolution visual generation. For instance, as illustrated in Fig. 2(a), using FP16 precision, SD-v1.5 fails to generate 2048-resolution images on A100, a GPU with 80GB of memory, due to out-of-memory errors, making higher resolutions or larger models even more problematic [^1] PyTorch 1.13 is adopted here for evaluation to reflect the theoretical complexity of various architectures. On higher versions of PyTorch, block-wise strategies are applied for memory efficient attention. However, the time efficiency is still a problem..
@@
The proposed generalized linear attention module is integrated into the architectures of SD, replacing the original self-attention layers, and the resultant model is termed as Linear-Complexity Diffusion Model, or LinFusion in short. By only training the linear attention modules for 50k iterations in a knowledge distillation framework, LinFusion achieves performance on par with or even superior to the original SD, while significantly reducing time and memory complexity, as shown in Fig. 2. Meanwhile, it delivers satisfactory zero-shot cross-resolution generation performance and can generate images at 16K resolution on a single GPU. It is also compatible with existing components and pipelines for SD, such as ControlNet Zhang et al. (2023), IP-Adapter Ye et al. (2023), DemoFusion Du et al. (2024), DistriFusion Li et al. (2024), etc, allowing users to achieve various purposes with the proposed LinFusion flexibly without any additional training cost. As shown in Fig. 1, extensive experiments on SD-v1.5, SD-v2.1, and SD-XL validate the effectiveness of the proposed model and method. Our contributions can be summarized as follows:
@@
  • Extensive experiments on SD demonstrate that the proposed LinFusion can achieve even better results than the original SD and exerts satisfactory zero-shot cross-resolution generation performance and compatibility with existing components and pipelines for SD. To the best of our knowledge, this is the first exploration of linear-complexity token mixers on the SD series model for text-to-image generation.
@@ 2.2 Overview
Technically, to derive a linear-complexity diffusion backbone, one simple solution is to replace all the self-attention blocks with Mamba2, as shown in Fig. 4(a). We apply bi-directional SSM to ensure that the current position can access information from subsequent positions. Moreover, the self-attention modules in Stable Diffusion do not incorporate gated operations Hochreiter & Schmidhuber (1997); Cho (2014) or RMS-Norm Zhang & Sennrich (2019) as used in Mamba2. As shown in Fig. 4(b), we remove these structures to maintain the consistency and result in a slight improvement in performance. In the following parts of this section, we delve into the issues of applying SSM, the core module in Mamba2, to diffusion models and accordingly introduce the key features in LinFusion: normalization and non-causality in Secs. 2.3 and 2.4 respectively. Finally, in Sec. 2.5, we provide the training objectives to optimize parameters in LinFusion modules.
@@ 2.4 Non-Causal Mamba
The proof can be found in the appendix. In practice, we adopt two MLPs to mimic the functionalities of $f^{\prime}$ and $g^{\prime}$. Combining with the normalization operations mentioned in Sec. 2.3, we derive an elegant structure shown in Fig. 4(d).
@@
The proof is provided in the appendix. From this point of view, the proposed structure can be deemed as a generalized linear attention and a non-causal form of recent linear-complexity sequential models, including Mamba2 Dao & Gu (2024), RWKV6 Peng et al. (2024), GLA Yang et al. (2023b), etc. In Tab. 1, we provide a summary of the parameterization in recent works for $A_{i}$.
@@ 3.3 Empirical Extensions
The proposed LinFusion is highly compatible with various components/pipelines for SD, such as ControlNet Zhang et al. (2023), IP-Adapter Ye et al. (2023), LoRA Hu et al. (2022), DemoFusion Du et al. (2024), DistriFusion Li et al. (2024), etc, without any further training or adaptation. We present some qualitative results in Fig. 5 and refer readers to the appendix for more results. The overall performance of LinFusion is comparable with the original SD.
@@
Ultrahigh-Resolution Generation. As discussed in Huang et al. (2024); He et al. (2024), directly applying diffusion models trained on low resolutions for higher-resolution generation can result in content distortion and duplication. A series of works are dedicated for higher-resolution image generation by leveraging off-the-shelf diffusion models Du et al. (2024); Lin et al. (2024a; b); Haji-Ali et al. (2024). However, limited by the quadratic-complexity self-attention, when applied for ultrahigh-resolution generation, existing approaches turn to patch-wise strategies to overcome the heavy computation burden Bar-Tal et al. (2023), which leads to inferior results, as shown in Settings A and B of Tab. 10.
Complementary to these methods, LinFusion addresses the computational overhead via generalized linear attention. As shown in Settings C and D of Tab. 10, LinFusion achieves 2$\sim$ 2$\times$ 2 × acceleration under $2048\times 2048$ resolution. Instead of going through full denoising steps in the original DemoFusion Du et al. (2024), tricks in SDEdit Meng et al. (2021) are additionally applied here so that the former 60% steps are skipped, which further enhances the efficiency without scarifying the quality. Please refer to the appendix for more analysis. Backed up by the linear-complexity LinFusion, such strategies enable ultrahigh-resolution generation up to 16K on a single GPU as shown in Fig. 1.
Distributed Parallel Inference. LinFusion is friendly for distributed parallel inference benefiting from its linear complexity, given that the communication cost is constant with respect to image resolution. Specifically, unlike the original DistriFusion Li et al. (2024) requiring transmitting all the key and value tokens for self-attention communication, the transmission in LinFusion is $g^{\prime}(X)^{\top}X\in\mathbb{R}^{c^{\prime}\times c}$, which is not related with the number of image tokens. In consequence, as shown in Tab. 3, LinFusion does not require NVLink hardware to achieve satisfactory acceleration. Please refer to the appendix for qualitative examples.
@@ 4 Conclusion
This paper introduces a diffusion backbone termed LinFusion for text-to-image generation with linear complexity in the number of pixels. At the heart of LinFusion lies in a generalized linear attention mechanism, distinguished by its normalization-aware and non-causal operations—key aspects overlooked by recent linear-complexity token mixers like Mamba, Mamba2, and GLA. We reveal theoretically that the proposed paradigm serves as a general low-rank approximation for the non-causal variants of recent models. Based on Stable Diffusion (SD), LinFusion modules after knowledge distillation can seamlessly replace self-attention layers in the original model, ensuring that LinFusion is highly compatible to existing components or pipelines for Stable Diffusion, like ControlNet, IP-Adapter, LoRA, DemoFusion, DistriFusion, etc, without any further training effort. Extensive experiments on SD-v1.5, SD-v2.1, and SD-XL demonstrate that the proposed model outperforms existing baselines and achieves performance on par with, or better than, the original SD with significantly reduced computational overhead. On a single GPU, it can accommodate image generation with resolutions up to 16K.
{floatrow}
[…]
Figure 9: Results of ControlNet on the original SD-v1.5 and LinFusion.
@@
[…]
Figure 10: Results of LinFusion on pipelines dedicated for high-resolution generation.
ID Setting FID($\downarrow$) CLIP-T($\uparrow$) Time($\downarrow$) (sec.)
A DemoFusion 70.01 0.343 61.36
B A - Patch 65.44 0.340 57.56
C B + SDEdit 65.15 0.344 26.98
D C + LinFusion 65.07 0.338 14.71
@@
{floatrow}
[…]
Table 3: Results of distributed parallel inference on a server with 8 RTX 4090 D GPUs. Benefiting from its linear complexity and constant communication cost among various patches, LinFusion is readily for distributed parallel inference with multiple GPUs. Compared with DistriFusion, it achieves more significant acceleration even without NVLink.
Appendix A Related Works
A.1 Efficient Diffusion Architectures
To mitigate the efficiency issue triggered by the quadratic time and memory complexity, a series of works, including DiS Fei et al. (2024a), DiM Teng et al. (2024), DiG Zhu et al. (2024), Diffusion-RWKV Fei et al. (2024b), DiffuSSM Yan et al. (2024), and Zigma Hu et al. (2024). These works have successfully adapted recent state space models like Mamba Gu & Dao (2023), RWKV Peng et al. (2023), or Linear Attention Katharopoulos et al. (2020) into diffusion architectures. However, these architectures maintain a causal restriction for diffusion tasks, processing input spatial tokens one by one, with generated tokens conditioned only on preceding tokens. In contrast, the diffusion task allows models to access all noisy tokens simultaneously, making the causal restriction unnecessary. To address this, we eliminate the causal restriction and introduce a non-causal token mixer specifically designed for the diffusion model.
A.2 Linear-Complexity Token Mixers
Appendix B Theoretical Proof
Proposition 1.
The proof is straightforward.
Proposition 2.
Given that $\tilde{A}=FG^{\top}$, $F,G\in\mathbb{R}^{n\times r}$, and $B,C\in\mathbb{R}^{n\times d^{\prime}}$, denoting $C_{i}=c(X_{i})$, $B_{i}=b(X_{i})$, $F_{i}=f(X_{i})$, and $G_{i}=g(X_{i})$, there exist corresponding functions $f^{\prime}$ and $g^{\prime}$ such that Eq. 4 of the main manuscript can be equivalently implemented as linear attention, expressed as $Y=f^{\prime}(X)g^{\prime}(X)^{\top}X$.
Proof.
Given existing conditions, we have:
$=[(c(X_{i})\otimes f(X_{i}))(b(X_{j})\otimes g(X_{j}))^{\top}]_{i,j},$
Proposition 3.
Proof.
Given existing conditions, we have:
Appendix C Additional Experiments
{floatrow}
[…]
Figure 13: LinFusion is complementary to pipelines dedicated to high-resolution generation like DemoFusion. To enhance the performance, instead of working patch-by-patch, we handle the image as a whole benefiting from the efficient LinFusion architecture. Moreover, we reveal that skipping part of the denoising steps in the high-resolution stage can further improve the efficiency without hurting the performance. The prompt is “An astronaut floating in space. Beautiful view of the stars and the universe in the background.”.
{floatrow}
[…]
Figure 14: LinFusion is complementary to pipelines for high-resolution generation like DistriFusion. Using constant communication cost among various GPU, it achieves highly comparable performance with single-GPU inference. The prompt is “Astronaut in a jungle, cold color palette, muted colors, detailed, 8k”.
{floatrow}
[…]
Figure 15: By default, LinFusion replaces all self-attention layers in a diffusion backbone. When applied to DiT-based structures like PixArt-Sigma, this configuration often struggles to generate smooth results. Leaving a small part of original self-attention layers unmoved, e.g., 25%, could largely alleviate this challenge. The prompts are “A photo of beautiful mountain with realistic sunset and blue lake, highly detailed, masterpiece” and “dog” respectively.
Ultrahigh-Resolution Generation. We present qualitative examples to illustrate the effectiveness of LinFusion on ultrahigh-resolution generation in Fig. 13. We build LinFusion upon DemoFusion Du et al. (2024), a pipeline dedicated to high-resolution generation. In the original implementation, for efficiency, DemoFusion handles a high-resolution image patch-by-patch and averages the outputs of overlapped areas Bar-Tal et al. (2023). However, we find that such a patch-wise treatment largely ignores the holistic text-image relationships. As shown in Fig. 13(DemoFusion), there are stars on the body of the astronaut. With an efficient architecture introduced by LinFusion, we do not have to conduct inference patch-by-patch. Instead, the whole image can be accommodated to a single GPU for denoising, which addresses the above limitation effectively as shown in Fig. 13(Full Steps).
Moreover, DemoFusion has to conduct full steps in the high-resolution denoising stage, which would introduce significant latency. Motivated from the insight in SDEdit Meng et al. (2021) that early denoising steps tend to take over the overall image layouts, we propose to skip some initial steps in the high-resolution stage, given that the overall image structures have been produced in the low-resolution stage. We find that it not only improves the efficiency but also makes the pipeline more robust to the turbulence on image layout in the high-resolution stage, as shown in Fig. 13(40% Steps).
Distributed Parallel Inference. We supplement qualitative results of distributed parallel inference by building LinFusion upon DistriFusion Li et al. (2024) in Fig. 14. Using constant communication cost among various GPUs, it achieves highly comparable performance with single-GPU inference. Unlike the original DistriFusion, LinFusion offers significant acceleration using multiple GPUs without the dependency on NVLink.
{floatrow}
[…]
Table 4: Performance of LinFusion built upon various models on the COCO benchmark.
Setting FID($\downarrow$) CLIP-T($\uparrow$)
SD-v1.5 12.86 0.321
w. LinFusion 12.57 0.323
SD-v2.1 12.84 0.333
w. LinFusion 13.84 0.329
SD-XL 14.74 0.340
w. LinFusion 15.72 0.338
PixArt-Sigma 26.32 0.334
w. 75% LinFusion 24.32 0.327
Results on More Architectures. We conduct experiments on a variety of diffusion architectures in this paper, including SD-v1.5, SD-v2.1 Rombach et al. (2022), SD-XL Podell et al. (2023), and PixArt-Sigma Chen et al. (2023). The former three adopt transformer-based UNet while the last one is based on DiT Peebles & Xie (2022), a pure-Transformer structure. Their quantitative results are listed in Tab. 4.
We find that on SD-v2.1 and SD-XL, LinFusion leads to slightly inferior results. We speculate that the reason lies in the training data used for LinFusion, which consists of only $\sim$160K relatively low-resolution samples, the majority of which are below $512\times 512$ resolution. Involving more high-quality samples can benefit the performance.
On PixArt-Sigma, we find that replacing all the self-attention layers in the DiT would result in unnatural results, as shown in Fig. 15. We speculate that the challenge arises because self-attention is the core and sole mechanism for managing token relationships in DiT. Replacing these layers entirely with LinFusion may create a significant divergence from the original architecture, leading to difficulties during training. As shown in Fig. 15, leaving a small part of original self-attention layers unchanged, e.g., 25%, could largely alleviate this challenge.
Appendix D Limitations
The motivation of LinFusion is to explore a linear-complexity diffusion architecture by experimentally replacing all the self-attention layers with the proposed generalized linear attention. This may not be the optimal configuration in practice. For example, it could be promising to explore hybrid structures and apply attention on deep features with a relatively smaller number of tokens but a large number of feature channels, which appears to be crucial for DiT-based architectures and could be a meaningful future direction.